where are the foci for an ellipse with vertices at (-5,0) and (5,0) and co-vertices at (0,3) and (0,-3) located?
Answer by lwsshak3(6519)

where are the foci for an ellipse with vertices at (-5,0) and (5,0) and co-vertices at (0,3) and (0,-3) located?
**
Standard form of equation for ellipse with horizontal major axis:
(x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k) being the (x,y) coordinates of the center.
..
Equation of given ellipse: x^2/25+y^2/9=1
For given ellipse:
Center: (0,0)
a=5
a^2=25
b=3
b^2=9
c^2=a^2-b^2=25-9=16
c=√16=4
Foci: (±c,0)=(±4,0)=(-4,0) and (4,0)